# Intensity-Efficient Allocations

### v5

@@ 4.3 Existence without Cardinality Assumptions
The argument proceeds by considering the possible ways in which an arbitrary profile $S\in\mathcal{\widehat{S}}$ when $n=3$ might generate a sequence of distinct Pareto-efficient allocations that are implicated in an intensity-dominance cycle. To this end, let $D$ be the intensity-dominance relation that is introduced in Definition 4. Suppose to the contrary that
@@
[[Note: in what follows we make repeated use—often without explicit reference—of the Pareto-efficiency implications (15)–(17), the skew-symmetry property $s_{l}(a,b)=-s_{l}(b,a)$ of canonical representations and, whenever exact values of the $s_{i}$ functions are asserted, of the lateral-consistency property (6) together with the assumption that every $s_{i}$ is canonical and strict (cf Observations 34).]]
@@
Since $w^{i}$ is Pareto efficient for $i\leq 4$, it follows from (15)–(17) that there are 4 subcases to consider:
@@
In view of (15)–(17), we can now consider the following 4 possible subcases:
Subcase 2-i. $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$. By (23) and (15)–(17), the former postulate implies $s_{3}(a,b)>0$. Since $s$ is canonical, this further implies $s_{3}(a,b)=1$, $s_{2}(a,b)=2$ and $s_{1}(a,b)=3$. This, canonicality of $s$ and $s_{1}(b,c)>s_{3}(b,c)>0$ together imply $s_{1}(b,c)=2$ and $s_{3}(b,c)=1=s_{3}(a,b)$, which contradicts Strictness.
@@
Subcase 3-i. $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$. The latter, together with (15)–(17), (25) and canonicality, implies $s_{3}(a,b)=1$. But canonicality, (15)–(17) and (17) also implies $s_{3}(a,b)=3$, which contradicts Strictness.
@@
Subcase 3-iii. $s_{2}(b,a)>s_{1}(b,a)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$. The first postulate, together with (26) and canonicality, implies $s_{2}(b,a)=3$, $s_{1}(b,a)=2$, $s_{3}(b,a)=1$. The second postulate and $s_{1}(b,a)=2$, together with Strictness, implies $s_{1}(b,c)=3$. This in turn implies $s_{1}(a,c)=1$ or $s_{1}(c,a)=1$. If the latter is true, then $s_{1}(b,c)>0$, $s_{1}(c,a)>0$ and (6), together with canonicality, implies $s_{1}(b,a)=3$, a contradiction. Hence, it must be that $s_{1}(a,c)=1$. We therefore have $s_{1}(b,a)=2$, $s_{1}(a,c)=1$ and, by (6) and canonicality, $s_{1}(b,c)=3$. From (15)–(17), (19), (25) and canonicality we also know that $s_{1}(b,c)>s_{3}(b,c)>s_{2}(b,c)>0$ implies $s_{3}(b,c)=2$ and $s_{2}(b,c)=1$. Thus, we have $s_{3}(b,a)>0$, $s_{3}(b,c)>0$ and, by (15)–(17) and $s_{1}(a,c)>0$, also $s_{3}(a,c)>0$. But $s_{3}(b,a)>0$, $s_{3}(a,c)>0$ together with (6) and canonicality implies $s_{3}(b,c)=3$, a contradiction.
@@
which, together with (18)–(19), imply
@@
Suppose first that $s_{2}(b,c)>0$. Then, by (29) and (15)–(17), $s_{1}(b,c)=3$, $s_{2}(b,c)=1$ and $s_{3}(b,c)=2$. From $s_{1}(b,c)=3$ and (6) we also get $s_{1}(b,a)>0$ and $s_{1}(a,c)>0$. Hence, by (18), $s_{2}(b,a)>s_{1}(b,a)>0$ and, by (27), $s_{2}(a,c)>s_{1}(a,c)>0$. These inequalities and (6) together imply $s_{2}(b,c)=3$, which is a contradiction.
Now suppose instead that $s_{2}(b,c)<0$, i.e. $s_{2}(c,b)>0$. It follows from (29) that $s_{2}(c,b)=3$, $s_{3}(c,b)=2$ and $s_{1}(c,b)=1$. Suppose $s_{1}(a,c)>0$. From (27) and (15)–(17), $s_{2}(a,c)>0$. Since $s_{2}(a,c)>0$ and $s_{3}(c,b)>0$, by (6) we get $s_{2}(a,b)>s_{2}(c,b)=3$, which is impossible. Hence, $s_{1}(c,a)>0$ holds instead and, from (27) and (15)–(17), $s_{1}(c,a)>s_{2}(c,a)>0$ is also true. Suppose $s_{2}(a,b)>0$ holds too. By (18), $s_{1}(a,b)>0$. By (6) and $s_{1}(c,a)>0$, $s_{1}(a,b)>0$ we get $s_{1}(c,b)=3$, a contradiction. Hence, $s_{2}(b,a)>0$ must be true instead and, by (18), $s_{2}(b,a)>s_{1}(b,a)>0$ also. So, we have $s_{2}(c,b)>0$ and $s_{2}(b,a)>0$, which, by (6), implies $s_{2}(c,a)>s_{(}c,b)=3$. This too is a contradiction.
@@
It has therefore been shown that $D$ is acyclic when $n=3$, establishing the existence of an intensity-efficient allocation in this case.
For the claimed potential non-existence when $n\geq 4$, consider the example strict intensity profile on $A=\{a,b,c,d\}$ whose canonical representation is shown in the table below:
$s_{i}(a,b)=$ $i=$ 1 2 3 4
6 $(a,d)$ $(a,d)$ $(a,d)$ $(a,d)$
5 $(b,d)$ $(a,c)$ $(a,c)$ $(b,d)$
4 $(a,c)$ $(b,d)$ $(b,d)$ $(a,c)$
3 $(b,c)$ $(b,c)$ $(b,c)$ $(b,c)$
2 $(a,b)$ $(c,d)$ $(c,d)$ $(a,b)$
1 $(c,d)$ $(a,b)$ $(a,b)$ $(c,d)$
We start by observing that $s_{1}=s_{4}$, $s_{2}=s_{3}$ and $a\succ_{i}b\succ_{i}c\succ_{i}d$, $i=1,\ldots,4$. This implies that all 24 possible allocations, defined and listed below, are Pareto efficient.
The following comparisons, whose validity can be readily established by the reader, demonstrate that for each $x^{i}$, $i=1,\ldots,24$, in this set there is a distinct $x^{j}$ that intensity-dominates $x^{i}$—a situation denoted by $x^{j}Dx^{i}$.
Therefore, no intensity-efficient allocation exists in this profile. ∎
The somewhat surprising fact that a plausible dominance concept may by cyclic and prevent an optimal entity to emerge invites an informal analogy between intensity-dominance cycles over allocations with at least 4 agents and Condorcet cycles over alternatives in pairwise majority-based preference aggregation with at least 3 agents (Condorcet, 1785). Importantly, however, although intensity-dominance cycles here may prevent refining the Pareto set, unlike the Condorcet social-welfare theoretic framework these do not lead to a “policy paralysis” problem because the Pareto set is always non-empty and, absent any distributively juster suggestions, one of them might be promoted by the social planner. That said, it is naturally of interest to understand better the conditions under which our proposed refinement of Pareto-efficiency is well-defined. This is the focus of the next subsection.
4.4 Existence Condition for General $n$: Homeo-Monotonicity
Definition 8.
Given a profile $S\in\mathcal{\widehat{S}}$, a set of agents $V^{\prime}\subseteq V$ is said to be:
  1. 1.
    Envy-free with respect to a set of items $A^{\prime}\subseteq A$ if each $v_{i}\in V^{\prime}$ has a distinct $\succ_{i}$-best item in $A$, and this item (denoted $b_{i}^{\prime}$) belongs to $A^{\prime}$.
  2. 2.
    Top-preference-coincident with respect to a set of items $A^{\prime}\subseteq A$ if $\succ_{i}^{A^{\prime}}=\succ_{j}^{A^{\prime}}$ for all $v_{i},v_{j}\in V^{\prime}$ and, when $A^{\prime}\subset A$, $a^{\prime}\in A^{\prime}$, $a\in A\setminus A^{\prime}$ holds, then $a^{\prime}\succ_{i}a$ for all $v_{i}\in V^{\prime}$.
An envy-free set $V^{\prime}$ with respect to $A^{\prime}$ consists of agents in $V^{\prime}$ who have distinct overall-best items in $A^{\prime}$. This terminology is motivated by the fact that, if $V^{\prime}\equiv V$ and $A^{\prime}\equiv A$, then each agent receiving their most preferred option results in an allocation that is Pareto-efficient and also envy-free (Varian, 1974; Moulin, 2004). A top-preference-coincident pair consists of agents whose preferences over $A^{\prime}$ are identical and also such that every item in $A^{\prime}$ is preferred to everything in $A\setminus A^{\prime}$, with no restrictions on preferences in the latter set.
Definition 9.
A set of agents $V^{\prime}\subseteq V$ that is top-preference-coincident with respect to a set of items $A^{\prime}\subseteq A$ at profile $S\in\mathcal{\widehat{S}}$ is monotonic with respect to $A^{\prime}$ if $|V^{\prime}|=|A^{\prime}|$ and
$\geq$ $\geq$ $\cdots$ $\geq$
$\geq$ $\geq$ $\cdots$ $\geq$
$\vdots$ $\vdots$ $\vdots$ $\vdots$ $\vdots$
$\geq$ $\geq$ $\cdots$ $\geq$
$s_{|V^{\prime}|}(a_{1},a_{2})$ >> $s_{|V^{\prime}|}(a_{2},a_{3})$ >> $\cdots$ >> $s_{|V^{\prime}|}(a_{|A^{\prime}|-1},a_{|A^{\prime}|})$ >> 0
Finally, a (strictly) monotonic such set is homeo-monotonic with respect to $A^{\prime}$ if
$s_{i}(a_{i},a_{j})$ $\geq$ $s_{j}(a_{i},a_{j})$ for all $1\leq i<j\leq|V^{\prime}|$.
Verbally, when $k$ agents have identical preferences over their top $k$ items, and agents can be (strictly) ordered in how much they prefer any two consecutively-ranked items in this subset [cf (strict) monotonicity], then this ordering cannot be violated in how much they prefer any non-consecutively ranked items in this subset (cf homeo-monotonicity).
Definition 10.
A profile $S\in\mathcal{\widehat{S}}$ is homeo-monotonic if, whenever $V$ and $A$ can be partitioned into subsets $V^{1},\ldots,V^{m}$ and $A^{1},\ldots,A^{m}$ where $V^{1}$ is envy-free with respect to $A^{1}$ and $V^{i>1}$ is monotonically top-preference-coincident with respect to $A^{i>1}$, then $V^{i>1}$ is homeo-monotonic with respect to $A^{i>1}$.
Definition 11.
Example 2.
The table below describes an effectively homeo-monotonic profile $S\in\mathcal{\widehat{S}}$ where $n=6$ and where the envy-free and top-preference-coincident sets that partition $V$ and $A$ are $(V^{1},A^{1})=(\{v_{1}\},\{a_{1}\})$, $(V^{2},A^{2})=(\{v_{2},v_{3},v_{4}\},\{a_{2},a_{3},a_{4}\})$, $(V^{3},A^{3})=(\{v_{5},v_{6}\},\{a_{5},a_{6}\})$:
$\overbrace{\hskip 7.0pt\text{$V^{1}$\hskip 7.0pt}}$ $\overbrace{\hskip 40.0pt\text{$V^{2}$}\hskip 40.0pt}$ $\overbrace{\hskip 20.0pt\text{$V^{3}$}\hskip 20.0pt}$
1 2 3 4 5 6
15 $(a_{1},a_{6})$ $(a_{2},a_{1})$ $(a_{2},a_{5})$ $(a_{2},a_{6})$ $(a_{5},a_{1})$ $(a_{5},a_{3})$
14 $(a_{2},a_{6})$ $(a_{2},a_{6})$ $(a_{3},a_{5})$ $(a_{3},a_{6})$ $(a_{5},a_{4})$ $(a_{5},a_{1})$
13 $(a_{1},a_{5})$ $(a_{2},a_{5})$ $(a_{2},a_{6})$ $(a_{4},a_{6})$ $(a_{5},a_{3})$ $(a_{5},a_{4})$
12 $(a_{3},a_{6})$ $(a_{2},a_{4})$ $(a_{3},a_{6})$ $(a_{2},a_{1})$ $(a_{5},a_{2})$ $(a_{5},a_{2})$
11 $(a_{2},a_{5})$ $(a_{3},a_{1})$ $(a_{2},a_{1})$ $(a_{3},a_{1})$ $(a_{6},a_{1})$ $(a_{6},a_{3})$
10 $(a_{1},a_{4})$ $(a_{3},a_{6})$ $(a_{3},a_{1})$ $(a_{4},a_{1})$ $(a_{6},a_{4})$ $(a_{6},a_{1})$
9 $(a_{4},a_{6})$ $(a_{3},a_{5})$ $(a_{4},a_{5})$ $(a_{5},a_{6})$ $(a_{5},a_{6})$ $(a_{5},a_{6})$
8 $(a_{3},a_{5})$ $(a_{2},a_{3})$ $(a_{2},a_{4})$ $(a_{2},a_{4})$ $(a_{6},a_{3})$ $(a_{6},a_{4})$
7 $(a_{2},a_{4})$ $(a_{3},a_{4})$ $(a_{2},a_{3})$ $(a_{2},a_{1})$ $(a_{2},a_{1})$ $(a_{2},a_{1})$
6 $(a_{1},a_{3})$ $(a_{4},a_{1})$ $(a_{3},a_{4})$ $(a_{4},a_{5})$ $(a_{2},a_{4})$ $(a_{4},a_{3})$
5 $(a_{5},a_{6})$ $(a_{4},a_{6})$ $(a_{4},a_{6})$ $(a_{5},a_{1})$ $(a_{3},a_{1})$ $(a_{2},a_{1})$
4 $(a_{4},a_{5})$ $(a_{5},a_{1})$ $(a_{1},a_{5})$ $(a_{1},a_{6})$ $(a_{6},a_{2})$ $(a_{6},a_{2})$
3 $(a_{3},a_{4})$ $(a_{4},a_{5})$ $(a_{4},a_{1})$ $(a_{2},a_{4})$ $(a_{2},a_{3})$ $(a_{2},a_{4})$
2 $(a_{2},a_{3})$ $(a_{5},a_{6})$ $(a_{1},a_{6})$ $(a_{2},a_{3})$ $(a_{3},a_{4})$ $(a_{4},a_{1})$
1 $(a_{1},a_{2})$ $(a_{6},a_{1})$ $(a_{6},a_{5})$ $(a_{3},a_{4})$ $(a_{4},a_{1})$ $(a_{1},a_{3})$
$\blacklozenge$
Remark 1.
If $V$ is envy-free with respect to $A$, then $S\in\mathcal{\widehat{S}}$ is effectively homeo-monotonic.
Clearly, any set of agents can be partitioned into groups according to whether distinct or identical alternatives are the most preferred within each group. The additional structure imposed by effectively homeo-monotonic profiles amounts to requiring that: (i) the preferences of agents with the same top-ranked item also coincide in how they rank their second-, third-, $\ldots$, $k$th-most preferred item for some $k\leq n$ that coincides with the number of agents in that group; (ii) the preferences of distinct such groups are sufficiently heterogeneous. While this condition is restrictive and one can easily construct profiles where intensity-efficient allocations exist but the condition is violated, the proof of the next result shows that it is helpful toward clarifying some important mechanisms for the existence of such allocations in the general case. Moreover, a weaker version of this condition that we consider later turns out to be necessary for such existence.
Theorem 2.
Proof.
@@ 1 $(a_{1},a_{2})$ $(a_{6},a_{1})$ $(a_{6},a_{5})$ $(a_{3},a_{4})$ $(a_{4},a_{1})$ $(a_{1},a_{3})$
Consider a profile $S\in\mathcal{\widehat{S}}$ with the postulated structure. Let $V^{1}:=\{v^{1}_{1},\ldots,v^{1}_{|V^{1}|}\}$ be the envy-free set with respect to $A^{1}:=\{a^{1}_{1},\ldots,a^{1}_{|A^{1}|}\}$, and let $V^{i>1}:=\{i_{1},\ldots,i_{|V^{i}|}\}$ be the $i$th top-preference-coincident set with respect to $A^{i>1}:=\{i_{1},\ldots,i_{|A^{i}|}\}$, $i=2,\ldots,m$. By the profile assumption, $|A^{i}|=|V^{i}|$ for all $i\leq m$, and the collections $(V^{i})_{i=1}^{m}$ and $(A^{i})_{i=1}^{m}$ partition the sets $V$ and $A$, respectively. Without loss of generality we may assume that $a^{1}_{j}$ is the top-ranked item of agent $v^{1}_{j}$, $j=1,\ldots,|A^{1}|$ (for the general case we apply a permutation over $\{1,\ldots,|A^{1}|\}$ on the $a^{1}$-subscripts). Recalling our earlier notation, this amounts to writing $a_{1}^{1}\equiv b_{1}^{1}$, $a_{2}^{1}\equiv b_{2}^{1}$, etc. Also without loss, for $A^{i>1}$, $V^{i>1}$ we may assume that $a^{i}_{1}\succ_{j}a^{i}_{2}\succ_{j}\cdots\succ_{j}a^{i}_{|A^{i}|}$ for all $v_{j}\in V^{i>1}$. (Note: the special cases where $V^{1}=\emptyset$, $V^{1}=V$ or $V^{i}=V$ for some $i>1$ are allowed by the postulated $S\in\mathcal{\widehat{S}}$, and so will they be in the proof.)
@@
For the “if” part, let $x$ be an arbitrary allocation that satisfies (i) and (ii). Assume to the contrary that $x$ is not Pareto efficient. Let $y$ be another allocation that Pareto-dominates $x$. This implies $s_{l}(y_{l},x_{l})\geq 0$ for every $l\leq n$, with strict inequality for some $l$. Let $V^{*}\subsetneq V$ consist of every agent $l$ for whom $s_{l}(y_{l},x_{l})=0$. By ($\mathbf{5}_{D}$), $s_{l}(y_{l},x_{l})=0\Leftrightarrow y_{l}=x_{l}$. Suppose $V^{*}=\emptyset$. From the above, this implies $s_{l}(y_{l},x_{l})>0$ for all $l\leq n$, with every agent assigned a different item under $x$ and $y$. If the envy-free set $V^{1}$ is non-empty, then this readily contradicts (i) for every $v_{l}\in V^{1}$. If $V^{1}=\emptyset$, then there is at least one $i>1$ such that the top-preference-coincident set $V^{i>1}$ is non-empty. Since $|V|=|A|=n$ holds by assumption, and recalling the notation that precedes the statement of Lemma 3, it follows from (ii) that for the specific agent $v_{1}^{i}\in V^{i>1}$ we must have $s_{1}^{i}(x_{1}^{i},x^{\prime}_{i})>0$ for all $x^{\prime}_{i}\in A\setminus\{x_{1}^{i}\}$, a contradiction. Now suppose $V^{*}\neq\emptyset$. Since $y$ Pareto-dominates $x$, the set $V^{**}:=V\setminus V^{*}$ that comprises every agent $l$ for whom $s_{l}(y_{l},x_{l})>0$ holds is non-empty. From the preceding argument we know that for every $v_{l}\in V^{**}$ it must be that $v_{l}\not\in V^{1}$ and $v_{l}\neq v_{1}^{i}\in V^{i>1}$, for all $i>1$. By $s_{l}(y_{l},x_{l})>0$ and $|A|=|V|$, there is some agent $v_{k}\neq v_{l}$ such that $y_{l}=x_{k}$. By ($\mathbf{5}_{D}$) and Pareto-dominance, $s_{k}(y_{k},x_{k})>0$. Since both $v_{k}$ and $v_{l}$ are improved upon by $y$, it follows that $v_{l}\in V^{i>1}$ and $v_{k}\in V^{j>1}$ for distinct top-preference-coincident sets $V^{i>1}$ and $V^{j>1}$ with respect to $A^{i>1}$ and $A^{j>1}$. By the homeo-monotonicity assumption, $A^{i}\cap A^{j}=\emptyset$. Since, by the definition of top-preference-coincident sets, $s_{l}(a,a^{\prime})>0$ and $s_{k}(b,b^{\prime})>0$ holds for all $a\in A^{i>1},a^{\prime}\in A\setminus A^{i>1}$ and $b\in A^{j>1},b^{\prime}\in A\setminus A^{j>1}$, and since, by (ii), it holds that $x_{l}\in A^{i>1}$ and $x_{k}\equiv y_{l}\in A^{j>1}$, the postulate $s_{l}(y_{l},x_{l})>0$ leads to a contradiction. Therefore, $x$ is Pareto-efficient.

### v6

@@ 6 Eliciting Ordinal Intensities and Improving Upon Pareto-Efficient Allocations
Proof of Theorem 1. The argument proceeds by considering the possible ways in which an arbitrary profile $S\in\mathcal{\widehat{S}}$ when $n=3$ might generate a sequence of distinct Pareto-efficient allocations that are implicated in an intensity-dominance cycle. To this end, let $D$ be the intensity-dominance relation that is introduced in Definition 4. Suppose to the contrary that
@@
[[Note: in what follows we make repeated use—often without explicit reference—of the Pareto-efficiency implications (22)(24), the skew-symmetry property $s_{l}(a,b)=-s_{l}(b,a)$ of canonical representations and, whenever exact values of the $s_{i}$ functions are asserted, of the lateral-consistency property (6) together with the assumption that every $s_{i}$ is canonical and strict (cf Observations 34).]]
@@
Since $w^{i}$ is Pareto efficient for $i\leq 4$, it follows from (22)(24) that there are 4 subcases to consider:
@@
In view of (22)(24), we can now consider the following 4 possible subcases:
Subcase 2-i. $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$. By (30) and (22)(24), the former postulate implies $s_{3}(a,b)>0$. Since $s$ is canonical, this further implies $s_{3}(a,b)=1$, $s_{2}(a,b)=2$ and $s_{1}(a,b)=3$. This, canonicality of $s$ and $s_{1}(b,c)>s_{3}(b,c)>0$ together imply $s_{1}(b,c)=2$ and $s_{3}(b,c)=1=s_{3}(a,b)$, which contradicts Strictness.
@@
Subcase 3-i. $s_{1}(a,b)>s_{2}(a,b)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$. The latter, together with (22)(24), (32) and canonicality, implies $s_{3}(a,b)=1$. But canonicality, (22)(24) and (24) also implies $s_{3}(a,b)=3$, which contradicts Strictness.
@@
Subcase 3-iii. $s_{2}(b,a)>s_{1}(b,a)>0$ and $s_{1}(b,c)>s_{3}(b,c)>0$. The first postulate, together with (33) and canonicality, implies $s_{2}(b,a)=3$, $s_{1}(b,a)=2$, $s_{3}(b,a)=1$. The second postulate and $s_{1}(b,a)=2$, together with Strictness, implies $s_{1}(b,c)=3$. This in turn implies $s_{1}(a,c)=1$ or $s_{1}(c,a)=1$. If the latter is true, then $s_{1}(b,c)>0$, $s_{1}(c,a)>0$ and (6), together with canonicality, implies $s_{1}(b,a)=3$, a contradiction. Hence, it must be that $s_{1}(a,c)=1$. We therefore have $s_{1}(b,a)=2$, $s_{1}(a,c)=1$ and, by (6) and canonicality, $s_{1}(b,c)=3$. From (22)(24), (26), (32) and canonicality we also know that $s_{1}(b,c)>s_{3}(b,c)>s_{2}(b,c)>0$ implies $s_{3}(b,c)=2$ and $s_{2}(b,c)=1$. Thus, we have $s_{3}(b,a)>0$, $s_{3}(b,c)>0$ and, by (22)(24) and $s_{1}(a,c)>0$, also $s_{3}(a,c)>0$. But $s_{3}(b,a)>0$, $s_{3}(a,c)>0$ together with (6) and canonicality implies $s_{3}(b,c)=3$, a contradiction.
@@
which, together with (25)(26), imply
@@
Suppose first that $s_{2}(b,c)>0$. Then, by (36) and (22)(24), $s_{1}(b,c)=3$, $s_{2}(b,c)=1$ and $s_{3}(b,c)=2$. From $s_{1}(b,c)=3$ and (6) we also get $s_{1}(b,a)>0$ and $s_{1}(a,c)>0$. Hence, by (25), $s_{2}(b,a)>s_{1}(b,a)>0$ and, by (34), $s_{2}(a,c)>s_{1}(a,c)>0$. These inequalities and (6) together imply $s_{2}(b,c)=3$, which is a contradiction.
Now suppose instead that $s_{2}(b,c)<0$, i.e. $s_{2}(c,b)>0$. It follows from (36) that $s_{2}(c,b)=3$, $s_{3}(c,b)=2$ and $s_{1}(c,b)=1$. Suppose $s_{1}(a,c)>0$. From (34) and (22)(24), $s_{2}(a,c)>0$. Since $s_{2}(a,c)>0$ and $s_{3}(c,b)>0$, by (6) we get $s_{2}(a,b)>s_{2}(c,b)=3$, which is impossible. Hence, $s_{1}(c,a)>0$ holds instead and, from (34) and (22)(24), $s_{1}(c,a)>s_{2}(c,a)>0$ is also true. Suppose $s_{2}(a,b)>0$ holds too. By (25), $s_{1}(a,b)>0$. By (6) and $s_{1}(c,a)>0$, $s_{1}(a,b)>0$ we get $s_{1}(c,b)=3$, a contradiction. Hence, $s_{2}(b,a)>0$ must be true instead and, by (25), $s_{2}(b,a)>s_{1}(b,a)>0$ also. So, we have $s_{2}(c,b)>0$ and $s_{2}(b,a)>0$, which, by (6), implies $s_{2}(c,a)>s_{(}c,b)=3$. This too is a contradiction.
@@
It has therefore been shown that $D$ is acyclic when $n=3$, establishing the existence of an intensity-efficient allocation in this case. The potential non-existence when $n\geq 4$ is proved in the main text. (cf Table 2). $\square$
@@
Proof of Theorem 2. Consider a profile $S\in\mathcal{\widehat{S}}$ with the postulated structure. Let $N^{1}:=\{N^{1}_{1},\ldots,N^{1}_{|N^{1}|}\}$ be the envy-free set with respect to $A^{1}:=\{a^{1}_{1},\ldots,a^{1}_{|A^{1}|}\}$, and let $N^{i>1}:=\{i_{1},\ldots,i_{|N^{i}|}\}$ be the $i$th top-preference-coincident set with respect to $A^{i>1}:=\{i_{1},\ldots,i_{|A^{i}|}\}$, $i=2,\ldots,m$. By the profile assumption, $|A^{i}|=|N^{i}|$ for all $i\leq m$, and the collections $(N^{i})_{i=1}^{m}$ and $(A^{i})_{i=1}^{m}$ partition the sets $N$ and $A$, respectively. Without loss of generality we may assume that $a^{1}_{j}$ is the top-ranked item of agent $N^{1}_{j}$, $j=1,\ldots,|A^{1}|$ (for the general case we apply a permutation over $\{1,\ldots,|A^{1}|\}$ on the $a^{1}$-subscripts). Recalling our earlier notation, this amounts to writing $a_{1}^{1}\equiv b_{1}^{1}$, $a_{2}^{1}\equiv b_{2}^{1}$, etc. Also without loss, for $A^{i>1}$, $N^{i>1}$ we may assume that $a^{i}_{1}\succ_{j}a^{i}_{2}\succ_{j}\cdots\succ_{j}a^{i}_{|A^{i}|}$ for all $v_{j}\in N^{i>1}$. (Note: the special cases where $N^{1}=\emptyset$, $N^{1}=N$ or $N^{i}=N$ for some $i>1$ are allowed by the postulated $S\in\mathcal{\widehat{S}}$, and so will they be in the proof.)
@@
For the “if” part, let $x$ be an arbitrary allocation that satisfies (i) and (ii). Assume to the contrary that $x$ is not Pareto efficient. Let $y$ be another allocation that Pareto-dominates $x$. This implies $s_{l}(y_{l},x_{l})\geq 0$ for every $l\leq n$, with strict inequality for some $l$. Let $N^{*}\subsetneq N$ consist of every agent $l$ for whom $s_{l}(y_{l},x_{l})=0$. By ($\mathbf{5}_{D}$), $s_{l}(y_{l},x_{l})=0\Leftrightarrow y_{l}=x_{l}$. Suppose $N^{*}=\emptyset$. From the above, this implies $s_{l}(y_{l},x_{l})>0$ for all $l\leq n$, with every agent assigned a different item under $x$ and $y$. If the envy-free set $N^{1}$ is non-empty, then this readily contradicts (i) for every $v_{l}\in N^{1}$. If $N^{1}=\emptyset$, then there is at least one $i>1$ such that the top-preference-coincident set $N^{i>1}$ is non-empty. Since $|V|=|A|=n$ holds by assumption, and recalling the notation that precedes the statement of Lemma 3, it follows from (ii) that for the specific agent $v_{1}^{i}\in N^{i>1}$ we must have $s_{1}^{i}(x_{1}^{i},x^{\prime}_{i})>0$ for all $x^{\prime}_{i}\in A\setminus\{x_{1}^{i}\}$, a contradiction. Now suppose $N^{*}\neq\emptyset$. Since $y$ Pareto-dominates $x$, the set $N^{**}:=N\setminus N^{*}$ that comprises every agent $l$ for whom $s_{l}(y_{l},x_{l})>0$ holds is non-empty. From the preceding argument we know that for every $v_{l}\in N^{**}$ it must be that $v_{l}\not\in N^{1}$ and $v_{l}\neq v_{1}^{i}\in N^{i>1}$, for all $i>1$. By $s_{l}(y_{l},x_{l})>0$ and $|A|=|N|$, there is some agent $v_{k}\neq v_{l}$ such that $y_{l}=x_{k}$. By ($\mathbf{5}_{D}$) and Pareto-dominance, $s_{k}(y_{k},x_{k})>0$. Since both $v_{k}$ and $v_{l}$ are improved upon by $y$, it follows that $v_{l}\in N^{i>1}$ and $v_{k}\in V^{j>1}$ for distinct top-preference-coincident sets $N^{i>1}$ and $N^{j>1}$ with respect to $A^{i>1}$ and $A^{j>1}$. By the homeo-monotonicity assumption, $A^{i}\cap A^{j}=\emptyset$. Since, by the definition of top-preference-coincident sets, $s_{l}(a,a^{\prime})>0$ and $s_{k}(b,b^{\prime})>0$ holds for all $a\in A^{i>1},a^{\prime}\in A\setminus A^{i>1}$ and $b\in A^{j>1},b^{\prime}\in A\setminus A^{j>1}$, and since, by (ii), it holds that $x_{l}\in A^{i>1}$ and $x_{k}\equiv y_{l}\in A^{j>1}$, the postulate $s_{l}(y_{l},x_{l})>0$ leads to a contradiction. Therefore, $x$ is Pareto-efficient.